5 An SRS of size 30 is taken from a populati Ље!ow, and in the data file srs30.dat. f size 100. The sample values are given 85266386 10 7 15 9 15 3 567 10 143 4 17 10 6 14 127 8 129 a What is the sampling weight for each unit in the sample? b Use the sampling weights to estimate the population total, . c Give a 95% CI for t. Does the fpc make a difference for this sample?
Answer #1

Soln

From the given sample,

Mean = 8.23

Std Deviation(s) = 3.99

n = 30

Std Error (SE) = 3.99/301/2 = 0.73

a)

Mean of the Weight = 8.23

b)

Population Mean = 8.23

c)

At alpha = 0.05,

ZCritical = 1.96

Hence

95% CI = Mean +/- ZCritical * SE = 8.23 +/- 1.96*0.73 = {6.8,15.04}